Supose the diameter of a circle is 50 feet long and has a chord of 48 feet long. What is the distance from the center of the circle to the chord.
Answer by khwang(438)


As the diagram below,chord AB = 48 ft,center O,
OA (radius) = 50/2 = 25,M is the mid point of AB. So, AM = 24.

A M B
---------
|
|
O
Since AM is perpendicular to AB, we have AO^2 = AM^2 + OM^2,
so OM^2 = AO^2 - AM^2 = 25^2 - 24^2 = 625 - 576 =49
Hence,the distance from the center of the circle to the chord is 7 ft
...Anwer